OpenStudy (itrymath):

A recent drop in sales has affected Stock D with the function g(x) = –6. Explain to Gordon how Stock D’s new price function, f(x) + g(x), will be created. Graph f(x) + g(x).

OpenStudy (itrymath):

@563blackghost

OpenStudy (itrymath):

@Zale101 @zepdrix

OpenStudy (itrymath):

omg thanks so much for coming.... i've been on this for literally 3 hours..

OpenStudy (itrymath):

okay so this question is related to number 1.... my function i created was f(x)=23(1.05)^x

563blackghost (563blackghost):

You are adding the two functions... \(\huge\bf{f(x)+g(x)=23(1.05)^{x}-6}\)

OpenStudy (itrymath):

no way..... are you serious! @563blackghost

OpenStudy (itrymath):

i admit i am the most retarded person ever.....

563blackghost (563blackghost):

All's good ^.^

563blackghost (563blackghost):

So you would graph it. https://www.calorful.com/graphing/calculator

OpenStudy (itrymath):

yeah i know.. used desmos xD

563blackghost (563blackghost):

Ooo ok well good :)

OpenStudy (itrymath):

thanks!!!